OpenStudy (delaneyparris):

find the range of each function for the domain {-4,-2,0,1.5,4}. f(x)=5x^2+4

OpenStudy (arbershabani97):

all you have to do is first use -4 in the equation for example f(-4)=5*(-4)^2+4 f(-4)=80+4 --> f(-4)=84 so the answers are {84,_,_,_,_} so you just continue and solve for each, so the answer for second one is f(-2)=.....